# Training a Scientific Reasoning Model for Chemistry

### v1

@@ 4.5 Distillation
To merge the specialist models into a final generalist model, we perform distillation via supervised fine tuning. This can also be seen as behavior cloning or expert iteration [91, 92].
Unlike previous work [17, 18, 19, 95], we do not rejection-sample model responses after training, but instead collect correct responses from the entire training run. These sequences are then filtered to remove those with low reasoning quality, as judged by an LLM and regex for non-english language (Section D.1). We further observed that some open-ended tasks are susceptible to answers with undesirable molecular substructures; we therefore reject such responses (Section D.2). Finally, if multiple responses remain for a given problem, only the two latest responses are kept.
@@ 4.6 Generalist RL
Having distilled all tasks into a single model, we perform a final GRPO phase across all tasks. An online curriculum is used (without seeding) to encourage learning. Unlike the task-specific phase, all accuracy rewards are binary, without any partial credit assigned. However, to disincentivize undesirable substructures arising during RL (after being rejected during distillation), we assign a molecule quality bonus reward during the last steps of this phase (Section D.2) As in the specialist phase, problem rewriting is enabled.
@@ 5 Results
Here we report the results of training Mistral-Small-24B-Instruct-2501 using the procedure outlined in Section 4. The seven specialist models were trained using 24-72 Nvidia H100 GPUs each, with a varying set of hyperparameters (detailed in Section E.1). A total of 186,010 sequences were collected from the specialist training runs for distillation. A single SFT epoch was sufficient for distillation, with a batch size of $64$ and learning rate of $1.9\times 10^{-5}$. The final all-task RL training phase was performed using 384 H100 GPUs, over 4 days; all hyperparameters are described in Section E.2.
@@ 5.1 Model Performance
An annotated model response is provided in Figure 4.
Figure 2: Per-task performance of our model compared to general-purpose LLMs. The human bar is an average of four contractors instructed not to use tools such as PubChem or internet searching, only the drawing tool ChemDraw. For multiple choice tasks, the “random” line accounts for varying numbers of options between problems.
@@ 5.2 Data Efficiency
Prior work has shown that reasoning models, when trained via RL with verifiable rewards (RLVR), can continually learn from as few as one training example [99]. In Section 5.1, we benchmark the performance of ether0 against other LLMs trained with and without reinforcement learning. In this section, we now investigate the data efficiency ofether0’s during both training and inference.
[…]
@@
First, we compare the performance of ether0 against a traditional model (i.e. not an LLM) trained with supervised learning. The Molecular Transformer (MT) [60] is a state-of-the-art model for chemical reaction prediction, trained on nearly 480,000 reactions from the USPTO dataset [100]. When trained on our dataset, containing approximately 60,000 reactions, ether0 outperforms MT, even when retraining MT on the same data (Figure 3A). Furthermore, ether0 outperforms MT on our held-out test set, achieving 70% accuracy after seeing just 46,000 training examples, compared to MT’s 64.1% when trained on the full USPTO dataset. Additionally, we retrained MT from scratch using our smaller dataset. The retrained versions of MT (denoted by MT) failed to exceed 30% accuracy, a threshold surpassed by ether0 after seeing only 10% of the available training data. This demonstrates that a reasoning model can achieve performance competitive with a dedicated traditional model given considerably less data.
@@
Second, we apply in-context learning (ICL) [2]) to evaluate the models’ ability to leverage additional data at inference time. ICL involves providing exemplar question-answer pairs directly in the prompt to guide the model’s response. In our setup, we construct ICL prompts from MCQs by selecting one of the distractors (i.e., incorrect options) from the original question and appending it as a labeled example. To maintain consistent random baselines between the one-shot and zero-shot versions, we remove the selected distractor from the set of choices in the actual question. Full details on the formatting and implementation of ICL are provided in Section F.4. Using this strategy, Figure 3B demonstrates a significant gain across MCQ tasks. Considering zero-shot performance, ether0 shows an overall performance of 50.1% in our test set, which is comparable to the 52.2% reached by ‘o1-2024-12-17‘. However, under one-shot prompting, ether0 surpasses all evaluated frontier models, highlighting its ability to generalize from minimal context. These results illustrate that our model, despite limited training data, can further increase performance and exceed the performance of frontier LLMs when appropriately guided at inference time.
@@
Figure 4: An annotated reasoning trace of the model correctly reasoning through an unseen structure elucidation task, for which o3, r1, Gemini 2.5-pro 05-07-25, and GPT 4.5 fail. The reasoning trace showcases examples of exploration, backtracking, and verification. The model doesn’t appear to know the real molecule name (azaleatin) and so calls it quercetin-C to indicate quercetin with an extra methyl group. Altogether this reasoning trace highlights both the strengths and limitations of ether0’s learned capabilities in the context of complex, multi-step chemical tasks.
@@ 5.3 Reasoning Performance and Behavior
In Figure 4, we annotate a representative completion of ether0 on a challenging open-answer task. The completion displays multiple lines of reasoning and verification, and additionally creates new words to help solve the problem, such as “Quercetin-C.” As judged by chemistry expert evaluation (Figure S6), the reasoning is generally coherent and proceeds logically from the question to the answer,
@@
Subsequently, we perform a more qualitative study of ether0’s reasoning. Recent work [25] suggests that the prevalence of “cognitive behaviors” (e.g. verification, backtracking) in a model’s reasoning is linked to its capacity to solve complex problems. To confirm this observation, we use a similar strategy to measure the frequency of such behaviors (behavior counts) over the course of model training (Section F.2).
@@ 6 Limitations
Although ether0 is trained on a variety of chemistry tasks, it may struggle to generalize when applied beyond its training distribution. For example, we do not expect strong performance on inorganic chemistry tasks, such as generating crystal structures, as the model was primarily trained on the SMILES strings of organic molecules. The intensive RL on these tasks also removed the general instruction-following and chat capabilities of Mistral-Small-24B-Instruct-2501 , including the capacity for multi-turn conversation. Finally, many small molecule design workflows involve heavy usage of tools. While the base model was trained with tool calling abilities, this was not part of ether0’s training. We leave it to future work to incorporate both chemistry reasoning and tool calling into a single model.
@@ 7 Conclusion
In this work, we show that reasoning models, which have demonstrated strong performance on tasks such as mathematics and programming, are also able to provide solutions to chemical reasoning questions that are often unsolvable by non-reasoning models. We introduce ether0, a 24B-parameter reasoning model trained on a curated dataset of challenging tasks focused on molecular design, completion, modification, and synthesis. We detail our training pipeline, which consists of several interleaved phases of reinforcement learning with verifiable rewards and behavior distillation. On a held-out evaluation set, ether0 significantly outperforms frontier LLMs, domain experts, and specialized models, particularly on open-answer tasks. We perform analysis of the model’s reasoning behavior, failure modes, and data efficiency, highlighting where reasoning helps and how it changes over training. Finally, we are releasing the model weights, benchmark data, and reward functions. We believe this work demonstrates strong potential for future work on reasoning models on scientific tasks.

### v2

@@ 4.3 Distillation
To merge the specialist models into a final generalist model, we perform another distillation via supervised fine tuning on the base Mistral-Small-24B-Instruct-2501 model. This can also be seen as behavior cloning or expert iteration 2017_Anthony; 2024_Havrilla.
Unlike previous work 2025_Bercovich; 2025_Akter; 2025_Blakeman; 2024_Abdin, we do not rejection-sample model responses after training, but instead collect correct responses from the entire training run. These sequences are then filtered to remove those with low reasoning quality, as judged by an LLM and regex for non-English language (Section D.1). We further observed that some open-ended tasks are susceptible to answers with undesirable molecular substructures; we therefore reject such responses (Section D.2). Finally, if multiple responses remain for a given problem, only the two latest responses are kept.
@@ 4.4 Generalist RL
Having distilled all tasks into a single model, we perform a combined GRPO phase across all tasks. An online curriculum is used (without seeding) to encourage learning. Unlike the task-specific phase, all accuracy rewards are binary, without any partial credit assigned. However, to disincentivize undesirable substructures arising during RL (after being rejected during distillation), we assign a molecule quality bonus reward during the last steps of this phase (Section D.2) As in the specialist phase, problem rewriting is enabled. We also run a safety alignment procedure described in Section E.3.
@@ 5 Results
Here we report the results of training Mistral-Small-24B-Instruct-2501 using the procedure outlined in Section 4. The seven specialist models were trained using 24-72 Nvidia H100 GPUs each, with a varying set of hyperparameters (detailed in Section E.1). A total of 186,010 sequences were collected from the specialist training runs for distillation. A single SFT epoch was sufficient for distillation, with a batch size of $64$ and learning rate of $1.9\times 10^{-5}$. The all-task RL training phase was performed using 384 H100 GPUs, over 4 days; all hyperparameters are described in Section E.2. The final safety alignment phase required 104 H100 GPUs (see Section E.3).
@@ 5.1 Model Performance
In Figure S2, we demonstrate that the our safety alignment procedure, which results in the ether0 refusing 80% of unsafe questions, does not meaningfully degrade capability on the measured tasks. An annotated model response is provided in Figure 4.
Figure 2: Per-task performance of our model compared to general-purpose LLMs. For multiple choice tasks, the “random” line accounts for varying numbers of options between problems. The human bar is an average of four chemists equipped with only the molecule drawing tool ChemDraw. Humans were not evaluated on receptor binding and scent tasks, as the structure-property relationship is mostly unknown, making these tasks essentially impossible without additional tools.
@@ 5.2 Data Efficiency
Prior work has suggested that training reasoning models via RL can be data-efficient 2025_WangSingleExample, although these results are not conclusive 2025_Chandak. In Section 5.1, we benchmark the performance of ether0 against other LLMs trained with and without reinforcement learning. In this section, we now investigate the data efficiency of ether0 during both training and inference.
Figure 3: Data efficiency analysis. (A) Comparison of ether0 to Molecular Transformer (MT) on reaction prediction: ether0 outperforms the published MT (dashed line) and shows higher data efficiency compared to retraining MT from scratch on our dataset ($\dagger$} - retrained). (B) Effect of in-context learning (ICL) on multiple-choice questions (MCQs).
@@
First, we compare ether0 to a traditional model (i.e., not an LLM) trained with supervised learning. The Molecular Transformer (MT) 2019_Schwaller is a state-of-the-art model for chemical reaction prediction, trained on nearly 480,000 USPTO reactions NIPS2017_USPTO. When trained on our dataset of  60,000 reactions, ether0 outperforms MT, even when MT is retrained on the same data (Figure 3A). On our held-out test set, ether0 achieves 70% accuracy after 46,000 examples, compared to MT’s 64.1% on the full USPTO dataset. We also retrained MT from scratch on our smaller dataset. The retrained versions of MT (denoted by MT) failed to exceed 30% accuracy, a threshold surpassed by ether0 after seeing only 10% of the available training data. This demonstrates that a reasoning model can achieve performance competitive with a dedicated traditional model given considerably less data.
@@
Second, we apply in-context learning (ICL) brown2020language) to evaluate the models’ ability to leverage additional data at inference time. ICL involves providing exemplar question-answer pairs directly in the prompt to guide the model’s response. In our setup, we construct ICL prompts from MCQs by selecting one of the distractors (i.e., incorrect options) from the original question and appending it as a labeled example. To maintain consistent random baselines between the one-shot and zero-shot versions, we remove the selected distractor from the set of choices in the actual question. Full details on the formatting and implementation of ICL are provided in Section F.4. Using this strategy, Figure 3B demonstrates a significant gain across MCQ tasks. Considering zero-shot performance, ether0 shows an overall performance of 50.1% in our test set, which is comparable to the 52.2% reached by ‘o1-2024-12-17‘. However, under one-shot prompting, ether0 surpasses all evaluated frontier models, highlighting its ability to generalize from minimal context. These results illustrate that our model, despite limited training data, can further increase performance and exceed the performance of frontier LLMs when appropriately guided at inference time.
@@
Figure 4: Annotated reasoning trace of the model correctly solving an unseen structure elucidation task, where o3, r1, Gemini 2.5-pro 05-07-25, and GPT-4.5 fail. The trace illustrates exploration, backtracking, and verification. The model does not know the real molecule name (azaleatin), referring to it as quercetin-C to indicate quercetin with an extra methyl group. Overall, this trace highlights both the strengths and limitations of ether0’s learned capabilities in complex, multi-step chemical tasks.
@@ 5.3 Reasoning Performance and Behavior
In Figure 4, we annotate a representative completion of ether0 on a challenging open-answer task. The completion displays multiple lines of reasoning and verification, and additionally creates new words to help solve the problem, such as “Quercetin-C.” As judged by chemistry expert evaluation (Figure S7), the reasoning is generally coherent and proceeds logically from question to answer.
@@
Subsequently, we perform a more qualitative study of ether0’s reasoning. Recent work gandhi2025cognitive suggests that the prevalence of “cognitive behaviors” (e.g. verification, backtracking) in a model’s reasoning is linked to its capacity to solve complex problems. To confirm this observation, we use a similar strategy to measure the frequency of such behaviors (behavior counts) over the course of model training (Section F.2).
@@ 6 Limitations
Although ether0 is trained on a variety of chemistry tasks, it can struggle to generalize beyond its training distribution. For example, we do not expect strong performance on inorganic chemistry, such as generating crystal structures, since the model was primarily trained on SMILES strings of organic molecules. The intensive RL training also reduced Mistral-Small-24B-Instruct-2501 ’s general instruction-following and chat capabilities, including multi-turn conversation. While many small-molecule design workflows rely heavily on tools, tool calling was not included in ether0’s training. In our evaluation, we used MT to validate predicted reactions, which may introduce limitations, and benchmarked against state-of-the-art LLMs, though other specialized non-LLM models could perform better on specific tasks. Future work could integrate chemistry reasoning and tool-calling into a single model.
@@ 7 Conclusion
In this work, we show that reasoning models, previously successful in mathematics and programming, can also solve chemical reasoning questions often unsolvable by non-reasoning models. We introduce ether0, a 24B-parameter reasoning model trained on a curated dataset of challenging tasks in molecular design, completion, modification, and synthesis. We detail our training pipeline, which consists of several interleaved phases of reinforcement learning with verifiable rewards and behavior distillation. On a held-out evaluation set, ether0 significantly outperforms frontier LLMs, domain experts, and specialized models, particularly on open-answer tasks. We analyze the model’s reasoning behavior, failure modes, and data efficiency, highlighting where reasoning helps and how it evolves during training. Finally, we release the model weights, benchmark data, and reward functions. We believe this work demonstrates strong potential for future work on reasoning models on scientific tasks.
@@ Acknowledgments
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We claim to be contributing a reasoning model trained for chemistry on various common chemistry tasks that outperforms specific deep-learning models on such tasks. This is extensively shown and supported by our results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section 6 discusses the limitations of our method, including the tasks our model does not perform too well.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A] .
  14. Justification: Our work does not include theoretical proofs. However, the needed background to understand the model training is included in this manuscript.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All the hyperparameters and training method are extensively described and included in this paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have publicly released a GitHub repository (https://github.com/Future-House/ether0) containing the reward functions and test data. All training data sources are publicly available, and we have also open-sourced the templates used to generate the question prompts.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Section 2 and SI Section C.2 are transparent on how the data was obtained and Section 4 discusses the training procedure in detail. The hyperparameters for every phase of the training pipeline are provided.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Training LLMs is very cost-intensive, and API calls to frontier LLMs are expensive. We considered a large evaluation set, but did not run replicates of each task.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The first paragraph of our “Results” section explicitly tells the compute resources used to train each step of our training pipeline.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our study does not have human subjects; all the data used is publicly available, and our model is being trained to mitigate safety and misuse concerns.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: We show that reasoning models are data efficient to learn chemistry-related tasks, which can be used to train a new model using our code for harmful purposes on a different dataset. This is not addressed in the main text. We briefly discuss potential applications of these models in drug discovery pipelines and the associated risks of misuse, but do not explore these topics in depth.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We reduce potential safety and misuse risks by incorporating a dedicated safety phase into the training pipeline, before releasing the nodel weights. We describe the full safety procedure in the SI.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Every source was thoroughly referenced, and a datasheet for the dataset is provided below.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our model has been open-sourced along with its model card, the reward function and dataset prompt templates are public and well documented in our GitHub repository (https://github.com/Future-House/ether0) and the dataset we use is derived from publicly available sources.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: Our study does not primarily involve crowdsourcing nor research with human subjects, but does pay a small number of human chemists to determine baseline accuracy on the problems being considered. Details on their compensation and instructions are included in the SI Section F.6.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our study does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Our study involves training LLMs, and LLMs were used in important steps of our method. All LLM used is well-described in Section 5.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.